OpenStudy (anonymous):

which of the following statements BEST describes the solution(s) of 1/3x^2 + 1/6x = 3 ?

OpenStudy (anonymous):

what are the statements? :P

OpenStudy (anonymous):

a) no solution b) 1 negative solution c) 1 positive and 1 negative solutions d) 2 positive solutions

OpenStudy (anonymous):

Okay. So the easiest thing to do is subtract the 3 across giving. (1/3)x^2+(1/6)x-3=0 I would multiply the whole thing by 6 to get rid of the fractions. 2x^2+x-18=0 using the quadratic formula you get: \[(-1\pm \sqrt{(1)^2-4(2)(18)})/(2*2) \]. This gives you a negative in the square root. So no solutions. Choice A. (in the real numbers that is :P)

OpenStudy (anonymous):

Thank you soo much dude (:

OpenStudy (anonymous):

No problem :)